OpenStudy (anonymous):

Given that Sinx=1/3, find tan2x.

OpenStudy (anonymous):

@burgandy U knw the formula for tan2x in terms of cosx and sinx?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[\tan(2x) =\frac{ 2tanx }{ 1-\tan^2x }\] here sinx is given so determine cosx first using \[\sin^2x +\cos^2x = 1\]

OpenStudy (anonymous):

remember tanx = sinx/cosx

OpenStudy (anonymous):

7/9?

OpenStudy (anonymous):

@312856MLP

OpenStudy (anonymous):

what is value of tanx u get?

OpenStudy (anonymous):

sinx = 3/5 that means opposite side of angle x is 3 and the hypotenuse is 5 so I don't think this is in a unit circle because unit circle has 1 as the radius or the hypotenuse. Since they gave sinx = 3/5 and I stated above, hypotenuse is 5, opposite is 3 so the other leg should be 4. Right triangle with sides 3, 4, 5. therefore cosx = 4/5 tanx = sinx/cosx = 3/4 By double angle formula tan(2x) = 2tanx/(1 - tan²u) tan(2x) = 2*(3/4)/[1 - (3/4)²] tan(2x) = (3/2)/(7/16) tan(2x) = 24/7 = 3 3/7

OpenStudy (anonymous):

apparently not the right one. :( would it be \[4\sqrt{2} /9\] ?

OpenStudy (anonymous):

do you get it

OpenStudy (anonymous):

@burgandy do u get it

OpenStudy (anonymous):

or /7?

OpenStudy (anonymous):

i m getting \[tanx = \frac{ 1 }{ 2\sqrt{2} }\] \[\tan2x = \frac{ \frac{ 2 }{ 2\sqrt{2} } }{ 1- {\frac{ 1 }{ 2\sqrt{2} }}^{2} }\]

OpenStudy (anonymous):

\7

OpenStudy (anonymous):

\[\frac{ \frac{ 1 }{ \sqrt{2} } }{ 1 - \frac{ 1 }{ 8 } }\] therefore \[\frac{ 8 }{ 7\sqrt{2} }\]

OpenStudy (anonymous):

But you can't have a radical on the bottom.

OpenStudy (anonymous):

it is 3 and 3/7

OpenStudy (anonymous):

ok \[\frac{ 8 }{ 7\sqrt{2} }*\frac{ \sqrt{2} }{ \sqrt{2} }\] \[\frac{ 8\sqrt{2} }{ 7*2 }\] \[\frac{ 4\sqrt{2} }{ 7 }\]

OpenStudy (anonymous):

is that ok?

OpenStudy (anonymous):

Yes! Thank you :)